OpenStudy (anonymous):

What is the equation for a circle with its center at the origin and a tangent whose equation is y=7?

OpenStudy (anonymous):

tangent to y=7 actually tells you the radius already.

OpenStudy (anonymous):

OKAY'

OpenStudy (anonymous):

just don't forget to square it (-:

OpenStudy (anonymous):

so it y=49

OpenStudy (anonymous):

(x - 5)2 + (y + 2)2 = 49 (x + 3)2 + (y - 3)2 = 49 x2 + y2 = 49

OpenStudy (anonymous):

a(x - 5)2 + (y + 2)2 = 49 b(x + 3)2 + (y - 3)2 = 49 cx2 + y2 = 49

OpenStudy (anonymous):

the equation of a circle looks like this \[ f(x,y)=(x-u)^2+(y-v)^2 = r^2 \] in this case (u,v) is the origin and therefore (0,0)

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thanks alot